OpenStudy (anonymous):

What is the 6th term in this geometric sequence? 480, 240, 120, 60, ...

OpenStudy (helder_edwin):

if u divide two consecutive terms u will get the ratio r of the sequence. and the n-th term would be \[\large a_n=a_{n-1}\cdot r \]

OpenStudy (anonymous):

i got 2 so then i plug it in?

OpenStudy (campbell_st):

well the terms are being halved.... and you have been given 4 terms... its easier to just halve 60 then havlve that answer for the 6th term

OpenStudy (helder_edwin):

no. the ratio is \[ \frac{120}{240}=\frac{1}{2} \]

OpenStudy (helder_edwin):

right. as @campbell_st just said.

OpenStudy (anonymous):

okay so then it would be 30 if its half, then 15, then 7.5?

OpenStudy (anonymous):

7.5 is the 6th term

OpenStudy (helder_edwin):

no. that is the 7th term.

OpenStudy (campbell_st):

30 is the 5th term so 15 is the 6th term

OpenStudy (anonymous):

oh right! counted ahead

OpenStudy (anonymous):

thanks guys!